# ”They’ve Stolen My GPL-Licensed Model!”: Toward Standardized and Transparent Model Licensing

### v1

@@ Appendix C Encoded Rules and Reasoning Logic
MGLicenseRule.ttl ¿ Llama2-derivative-rule @prefix mg: <http://~/rdf/terms#> . mg:Llama2-derivative-rule a mg:Rule ; mg:hasOutputDef mg:derivative ; mg:targetActionType mg:Amalgamate, mg:Combine, mg:Modify, mg:Train, mg:Embed, mg:Distill ; mg:targetInputWorkForm mg:weights, mg:exe ; mg:targetOutputWorkForm mg:weights, mg:exe ; mg:relicense mg:any-license ; mg:hasPublishRestriction mg:include_license_restriction ; mg:hasUseRestriction mg:llama2_exclusive_use_restriction, mg:use_behavior_restriction ; mg:allowSharing true .
@@
ruling.n3 ¿ License Determination Logic (snippet) @prefix log: <http://www.w3.org/2000/10/swap/log#> . @prefix list: <http://www.w3.org/2000/10/swap/list#> . @prefix mg: <http://~/rdf/terms#> . { # CASE-3: There have multiple rulings require the output work’s license # should be compatible, and a compatible solution exists. ?outw a mg:Work . _:x log:notIncludes { ?outw mg:hasLicense ?li } . # If all relied wroks have a license, we can determind the license of this work. ( {?outw mg:hasReliedwork ?relw} {?relw mg:hasLicense ?li} ) log:forAllIn _:t . # There is no ruling that NOT allow relicesnse (For exclude CASE-2). ( {?outw!mg:hasRuling!mg:hasRule mg:relicense ?reli } {?reli log:notEqualTo mg:none-license} ) log:forAllIn _:t . # Collect all compatible-relicensable licenses into a list. ( ?li { ?outw mg:hasRuling ?rling . ?rling!mg:hasRule mg:relicense ?rule . ?rule log:equalTo mg:compatible-license . # Compatible ?rling mg:hasLicense ?li . ?li mg:hasCompatibleLicense ?clist . } ?li_list ) log:collectAllIn _:t . # (CASE-1 will yield an empty list here) . ?li_list list:first ?li_1st . ?li_1st mg:hasCompatibleLicense ?compat_list_1st . _:x log:includes { ?cli list:in ?compat_list_1st . ({?li_list!list:member mg:hasCompatibleLicense ?compat_list_other} {?cli list:in ?compat_list_other} ) log:forAllIn _:t . } . } => { ?outw mg:hasLicense ?cli . } .
@@
analysis_granting.n3 ¿ Right Reserved Error { # Error [Right Reserved]. License reserves the right for your action. ?req a mg:Request . ?req!mg:grant mg:hasUsage ?req_usage . # There may be multiple mg:Usage. ?req mg:targetAction ?a . ?req mg:targetWork ?outw . ?req <- mg:hasRequest ?inw . ?inw mg:hasLicense ?li . # The checking target work must have a license. # Collect all reserved rights according to the license. ( ?r { ?li!mg:reserve mg:hasUsage ?r . } ?reserved_list ) log:collectAllIn _:x . ?req_usage list:in ?reserved_list . ( ?a ?inw ?outw ?req_usage ) log:skolem ?geniri . # Keep unique (”**Error**[RightReserved] ?req_usage isreservedby ?li licensefor ?a actionon ?inw toproduce ?outw) string:concatenation ?content . } => { ?geniri a mg:Error ; mg:reportBy ?a ; mg:reportType ”RightReserved” ; mg:content ?content . } .
Appendix D Supplementary of MG Licenses
Table 6 and Table 7 present the Model Sheet included in the Attachment of our proposed MG Licenses. For instance, the Model Sheet for MG-BY-NC indicates that this license is revocable, prohibits commercial use of the model, its output, associated code, documentation, and derivatives, and does not require disclosure of source weights or impose responsible AI restrictions. In comparison, MG-BY-OS has fewer restrictions, with its primary requirement being the disclosure of model weights.
Table 6. Model Sheet of MG-BY-NC.
Use & Modify Sublicensing
Create Derivatives Irrevocable
Share Verbatim Copy Trademark Use
Share Derivatives Commercial Use of Model
Share Output Commercial Use of Output
Patent Use Commercial Use of Derivatives
Copyright Use Commercial Use of Code & Docs
Retain Original Attribution Disclose Source
Retain Original License Responsible AI Restrictions
Retain All Notices
Disclaimer of Warranty
Limitation of Liability
Table 7. Model Sheet of MG-BY-OS.
Use & Modify Trademark Use
Create Derivatives Responsible AI Restrictions
Share Verbatim Copy
Share Derivatives
Share Output
Patent Use
Copyright Use
Commercial Use of Model
Commercial Use of Output
Commercial Use of Derivatives
Commercial Use of Code & Docs
Retain Original Attribution
Retain Original License
@@ Appendix D Supplementary of MG Licenses
[…]
@@ Appendix E RDF Workflow Visualization
Figure 6 illustrates the visualization of the workflow constructed by the MG Analyzer, as presented in the OSS licenses example in Section 5, Setting (i), which involves the publishing of Data D. This figure was generated using RDF Grapher, provided by the Linked Data Finland platform (Hyvönen et al., 2014). Note that, for clarity, only a subset of the properties is displayed.
[…]

### v2

@@ Appendix B Encoded Rules and Reasoning Logic
rules_ruling.n3 ¿ License Determination Logic (snippet) @prefix log: <http://www.w3.org/2000/10/swap/log#> . @prefix list: <http://www.w3.org/2000/10/swap/list#> . @prefix mg: <http://~/rdf/terms#> . { # CASE-3: There have been multiple rulings requiring the output work’s license # to be compatible, and these rulings have the same (compatible) license. ?outw a mg:Work . _:x log:notIncludes { ?outw mg:hasLicense ?li } . # If all relied works have a license, we can determine the license of this work ({?outw mg:hasReliedwork ?relw } { ?relw mg:hasLicense ?li}) log:forAllIn _:t . # There is no ruling that does NOT allow relicensing (to exclude CASE-2) ( {?outw!mg:hasRuling!mg:hasRule mg:relicense ?reli } {?reli log:notEqualTo mg:none-license} ) log:forAllIn _:t . # Collect all compatible relicensable licenses into a list ( ?li { ?outw mg:hasRuling ?rling . ?rling!mg:hasRule mg:relicense ?rule . ?rule log:equalTo mg:compatible-license . # Compatible ?rling mg:hasLicense ?li . # This ruling should bind a license which provides a compatible list ?li mg:hasCompatibleLicense ?clist . } ?li_list ) log:collectAllIn _:t . # (CASE-1 will yield a empty list here). ?li_list list:first ?li_1st . ?li_1st mg:hasCompatibleLicense ?compat_list_1st . # Check whether there exists a compatible license from the first list # that also appears in all other (including itself) compatible lists. # Return true if there is a license included in all compatible list. _:x log:includes { ?cli list:in ?compat_list_1st . ( {?li_list!list:member mg:hasCompatibleLicense ?compat_list_other} {?cli list:in ?compat_list_other} ) log:forAllIn _:t . } . } => { ?outw mg:hasLicense ?cli } .
@@
rules_analysis_granting.n3 ¿ Right Reserved Error { # Error [Right Reserved]. License reserves the right for your action. ?req a mg:Request . ?req!mg:grant mg:hasUsage ?req_usage . # There may be multiple mg:Usage. ?req mg:targetAction ?a . ?req mg:targetWork ?outw . ?req <- mg:hasRequest ?inw . ?inw mg:hasLicense ?li . # The target work must have a license. # Collect all reserved rights according to the license. ( ?r { ?li!mg:reserve mg:hasUsage ?r . } ?reserved_list ) log:collectAllIn _:x . ?req_usage list:in ?reserved_list . ( ?a ?inw ?outw ?req_usage ) log:skolem ?geniri . # Keep unique (”**Error**[RightReserved] ?req_usage isreservedby ?li licensefor ?a actionon ?inw toproduce ?outw) string:concatenation ?content . } => { ?geniri a mg:Error ; mg:reportBy ?a ; mg:reportType ”RightReserved” ; mg:content ?content . } .
Table 4. Supported Actions in MG Analyzer Following Rule Alignment. The symbols == and $\approx$ indicate that the Type/Form of output work and input work are the same and may differ, respectively. The corresponding OSS, Free-content, and Model license terms for each action are listed.
Action Type Form Composition Terms
Copy == == Output and input are exactly same. Copy Duplicate
Modify == == Output includes a significant portion of input and can be reverted. Modify Fine-tune
Amalgamate == == Output includes portions of input but cannot be reverted. Modify Remix Fusion
Train == == Output has the same structure as input and may contain a negligible portion of it. Alter Adapt Train
Synthetic
Distill == == Output does not contain any portion of input but performs similarly to it. Distill Transfer
Extract
Embed == == The output contains no portion of the input, but a mapping converts input to output. Translate Transform
Publish == $\approx$ Output is the same as the input but may have a different form. Redistribute Perform Display Disseminate
Appendix C Supplementary of MG Licenses
@@
Appendix D Supplementary License Analysis Results: Complex In-the-Wild Workflow
Figure 6. Analysis Results for an In-the-Wild Workflow based on SmolLM. Detected Issues Are Indicated in the Lower-Right Corner of Each Reported Work.
Figure 8. RDF Graph Visualization of the Compositional Dependency Reasoned by MG Analyzer for the SmolLM Example Workflow in Appendix D.
While MG Analyzer is primarily designed for pre-deployment compliance analysis of ML workflows, it can also be applied to in-the-wild workflows, provided that licenses for intermediate assets are pre-assigned. In this section, we present the analysis results of a SmolLM-based workflow (Allal et al., 2024). Its reasoned compositional dependencies are shown in Figure 8, and the production chain is illustrated in Figure 6. This complex workflow involves the generation of synthetic data for training and finetuning, as well as dataset combination. Although most assets have been relicensed under permissive licenses, our analyzer still identifies six notices and six warnings. Note that, the Stack v2 (Lozhkov et al., 2024) is a crowdsourced dataset containing code under permissive or public domain licenses. In our analysis, we treat its license as Unlicense. For demonstration purposes, we assume this workflow is constructed and published in a single step, rather than through repeated construct-then-publish cycles. The most commonly reported warning is License Type Mismatch, which can lead to increased ambiguity in license interpretation and may raise the risk of misusing licensed works. Another warning is triggered by Llama 3.1 and ODC-By, both categorized as Possibly Revocable Licenses. Warning W7 concerns the use of the everydata conversations dataset, which must comply with the Use Behavior restrictions specified in the Llama 3.1 license. The six notices are triggered by finetuning on models licensed under Apache-2.0, which requires including the license file, providing proper notices, and stating any changes when redistributing derivative works.
It is worth emphasizing that legal noncompliance issues can become increasingly difficult to address as the reuse chain evolves. For example, Use Behavior restrictions are often obscured through substream model reuse over multiple generations. The model cards (Mitchell et al., 2019) of Everyday Conversations and SmolLM-360M-Instruct, for instance, do not mention such constraints, posing a potential risk of unintended license violations. Moreover, because detailed ML workflow metadata is not systematically recorded or standardized in current practice, it is challenging to obtain complete reproduction information for accurate analysis. This may compromise the comprehensiveness of our results. Please refer to the Appendix A for a discussion of our work’s limitations.
@@ Appendix E RDF Workflow Visualization
Figure 7 illustrates the workflow constructed by the MG Analyzer, as presented in the OSS licenses example (Section 5, Setting (i)) involving the publishing of Data D. Figure 8 provides a visualization for the case discussed in Appendix D. Both figures were generated using RDF Grapher, a tool provided by the Linked Data Finland platform (Hyvönen et al., 2014). Note that, for clarity, only a subset of the properties is displayed.